OpenStudy (anonymous):

??? What is the equation of the line passing through (2, –2) and perpendicular to the line y = –x + 3 in slope-intercept form?

OpenStudy (anonymous):

Can you tell me what is the slope of perpendicular line given: Hint : Tell me the coefficient of x..

OpenStudy (anonymous):

??

OpenStudy (anonymous):

2

OpenStudy (anonymous):

In y = -x + 3 tell me the coefficient of x here..

OpenStudy (anonymous):

-x?

OpenStudy (anonymous):

It is the number attached to the x.. If I say -2x then -2 is the coefficient of x.. Likewise what is the coefficient of x in y = -x + 3??

OpenStudy (anonymous):

-2?

OpenStudy (anonymous):

How -2?? You are given with y = -x + 3 Here, you can write it as: y = (-1)x + 3 Now can you tell me the coefficient or not??

OpenStudy (anonymous):

-1

OpenStudy (anonymous):

Yes.. This is the slope of the given line.. So, firstly we have to find the slope of perpendicular line: \[\large m_1 m_2 = -1\] Here, \(m_1 = -1\), can you find \(m_2 \) from here??/

OpenStudy (anonymous):

-2 ?? o.O?

OpenStudy (anonymous):

you will get \(m_2 = 1\) So put in the equation: \[\large y - y_1 = m(x - x_1)\] \[y + 2 = 1(x - 2) \implies y + 2 = x - 2 \implies x - y = 4\]

OpenStudy (anonymous):

oh so that would be y=x+4 right?